Hi-Tek Manufacturing, Inc., makes two types of industrial component parts—the B300 and the T500. An absorption costing income statement for the most recent period is shown:

Hi-Tek Manufacturing Inc.
Income Statement
Sales $ 1,705,200
Cost of goods sold 1,242,098
Gross margin 463,102
Selling and administrative expenses 610,000
Net operating loss $ (146,898 )

Hi-Tek produced and sold 60,300 units of B300 at a price of $20 per unit and 12,800 units of T500 at a price of $39 per unit. The company’s traditional cost system allocates manufacturing overhead to products using a plantwide overhead rate and direct labor dollars as the allocation base. Additional information relating to the company’s two product lines is shown below:

B300 T500 Total
Direct materials $ 400,900 $ 162,100 $ 563,000
Direct labor $ 120,500 $ 42,600 163,100
Manufacturing overhead 515,998
Cost of goods sold $ 1,242,098

The company has created an activity-based costing system to evaluate the profitability of its products. Hi-Tek’s ABC implementation team concluded that $59,000 and $106,000 of the company’s advertising expenses could be directly traced to B300 and T500, respectively. The remainder of the selling and administrative expenses was organization-sustaining in nature. The ABC team also distributed the company’s manufacturing overhead to four activities as shown below:

Manufacturing
Overhead
Activity
Activity Cost Pool (and Activity Measure) B300 T500 Total
Machining (machine-hours) $ 204,618 90,400 62,300 152,700
Setups (setup hours) 150,880 78 290 368
Product-sustaining (number of products) 100,000 1 1 2
Other (organization-sustaining costs) 60,500 NA NA NA
Total manufacturing overhead cost $ 515,998

Required:

1. Compute the product margins for the B300 and T500 under the company’s traditional costing system.

2. Compute the product margins for B300 and T500 under the activity-based costing system.

3. Prepare a quantitative comparison of the traditional and activity-based cost assignments.
Answer #1

Part A

Predetermined overhead rate = Estimated total manufacturing overhead cost
Estimated total direct labor dollars = 515998/163100 = $3.16 per direct labor dollar

B300

T500

Total

Sales revenue

1206000

499200

1705200

Direct materials

400900

162100

563000

Direct labor

120500

42600

163100

Manufacturing overhead applied (@ $3.16 per direct labor-dollar)

380780

134616

515396

Total manufacturing cost

902180

339316

1241496

Product margin

303820

159884

463704

60300*20 =1206000

12800*39 =499200

Part B

Activity Cost Pools

total cost

total activity

activity rate

machining

204618

152700

1.34

set ups

150880

368

410

product sustaining

100000

2

50000

B300

T500

Total

Sales revenue

1206000

499200

1705200

Direct materials

400900

162100

563000

Direct labor

120500

42600

163100

machining

121136

83482

204618

set ups

31980

118900

150880

product sustaining

50000

50000

100000

Advertising expense

59000

106000

165000

Total manufacturing cost

783516

563082

1346598

Product margin

422484

(63882)

358602

Part C

The quantitative comparison is as follows:

Traditional Cost System

B300

%

T500

%

Total

Direct materials

400900

71.21%

162100

28.79%

563000

Direct labor

120500

73.88%

42600

26.12%

163100

Manufacturing overhead applied (@ $3.24 per direct labor-dollar)

380780

73.88%

134616

26.12%

515396

Total cost assigned to products

902180

339316

1241496

Selling and administrative expense

610000

Total cost

$1851496

Activity-Based Costing System

B300

%

T500

%

Total

Direct costs:

Direct materials

400900

71.21%

162100

28.79%

563000

Direct labor

120500

73.88%

42600

26.12%

163100

Advertising expense

59000

35.76%

106000

64.24%

165000

Indirect costs

machining

121136

59.20%

83482

40.80%

204618

set ups

31980

21.20%

118900

78.80%

150880

product sustaining

50000

50.00%

50000

50.00%

100000

Total cost assigned to products

783516

563082

1346598

Costs not assigned to products:

Selling and administrative

445000

other

60500

Total cost

$1852098

610000-165000 = 445000